5. A recent study shows that the annual cost of maintaining a building in Accra averages
$200 with a variance of $260. If a tax of 30% is introduced on all items associated with the maintenance of building i.e. everything is made 30% more expensive. Calculate the variance of the annual cost of maintaining a building in Accra.

6. Suppose that an airline accepted 12 reservations for a commuter plane with 10 seats.
They know that 7 reservations went to regular commuters who will show up for sure. The other 5 passengers will show up with a 50% chance, independently of each other.
Find the probability that the flight will be overbooked.

7. Suppose that 3% of computer chips produced by a certain machine are defective. The
chips are put into packages of 20 chips for distribution to retailers. What is the probability that a randomly selected package of chips will contain at least 2 defective
chips.
1
Expert's answer
2020-06-03T17:56:04-0400

5) Let X represents the annual cost of maintaining a building in Accra, then, E(x)=200, V(x)=260,Let Y represents the annual cost of maintaining a building in Accra after adding a tax of 30%,Y=1.3X

V(Y)=V(1.3X)=1.32V(x)=1.32(260)=439.4

6) Find the probability that the flight will be overbooked.

P(X=4)+P(X=5)= C5,4(0.5)4(1−0.5)5−4+C5,5(0.5)5(1−0.5)5−5=(5+1)(0.5)5=0.1875

7) n=20,p=0.03

P(x≥2)=1-P(x=0)-P(x=1)

P(x=0)=C20,0*0.030*(1-0.03)20-0=0.5438

P(x=1)=C20,1*0.031*0.9719=0.3363

P(x≥2)=1-0.5438-0.3363=0.12